. 1. How much mass of ammonia is formed when 3.12g of nitrogen is reacted with hydrogen
1
Expert's answer
2020-07-22T05:41:34-0400

Nitrogen and hydrogen react to form ammonia according to the equation:

N2 + 3H2 = 2NH3

In the equation, nitrogen and hydrogen reacts in the ratio 1:3

let consider reaction at atomic level by calculating the reacting number of moles

the number of moles is calculated from the formula:

moles = mass/molar mass

moles of nitrogen = 3.12/14, molar mass of nitrogen = 14

moles of nitrogen = 0.22286

The ratio of nitrogen and ammonia 1:2

If ratio 1 has 0.22286 moles, thus ratio 2 has 2*0.22286 = 0.44571 moles of ammonia formed

mass of ammonia = moles*molar mass

molar mass of ammonia= 14+3 = 17

Hence, mass of ammonia = 7.57707, which is approximately 7.56g, ammonia formed